# Universal Definitions of the Roman Factorial: Introduction to Foundational Functions and the Generalization Process

### v2

@@ 3.3 The function $\theta(n)$
$\theta(n)=\begin{cases}+1\thinspace,\quad n\geq 0\\ -1\thinspace,\quad n<0\end{cases},\enspace n\in\mathbb{R}\thinspace.$ (Eq. 3.3.1)
@@
$\theta(n)=\frac{\delta(n)}{\left|\delta(n)\right|}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.3.2)
@@ 3.4 The function $\xi(n)$
$\xi(n)=\begin{cases}1\thinspace,\quad n\geq 0\\ 0\thinspace,\quad n<0\end{cases},\enspace n\in\mathbb{R}\thinspace.$ (Eq. 3.4.1)
Let’s construct $\xi(n)$ using $\theta(n)$ step by step. To begin with, we note that the output pattern of $\theta(n)$ is $[-1,1,1]$. Adding 1 to $\theta(n)$ essentially adds to each output, obtaining $[0,2,2]$. Dividing $\bigl{(}1+\theta(n)\bigr{)}$ by 2 yields the desired pattern $[0,1,1]$. Thus, we arrive at the universal definition of $\xi(n)$:
$\xi(n)=\frac{1+\theta(n)}{2}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.4.2)
@@
This function has similarities to true/false states in Boolean logic, which is commonly used in programming. Specifically, $\xi(n\geq 0)=1$ and $\xi(n<0)=0$ represent the true/false statements of the condition $n\geq 0$. Furthermore, it’s noteworthy that the most subsequent F.F. can also be expressed in Boolean algebra.
@@ 3.5 The function $\xi^{\prime}(n)$
The function $\xi^{\prime}(n)$, marked with (${}^{\prime}$), bears resemblance to $\xi(n)$, although it is not its derivative. In this paper, the (${}^{\prime}$) symbol signifies similarity between functions rather than indicating a derivative. Notably, the F.F. are devoid of derivatives since they are constructed upon the floor function, which itself lacks a derivative at integers. Derivatives will not be discussed within the scope of this work [^9] It should be mentioned again that the names of the F.F. are arbitrarily assigned Greek or Latin letters and have no relation to other functions with the same name. The naming process was conducted without any specific goal or consideration of potential overlaps with pre-existing functions..
@@
$\xi^{\prime}(n)=\frac{1-\theta(n)}{2}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.5.1)
@@
$\xi^{\prime}(n)=1-\xi(n)\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.5.2)
In Boolean algebra, $\xi^{\prime}(n)$ yields the true/false states of the condition $n<0\thinspace$, thereby exhibiting an output pattern of $[1,0,0]$.
@@ 3.6 The function $\eta(n)$
$\eta(n)=\begin{cases}\thinspace 1\thinspace,&n\geq 0\\ \thinspace 1\thinspace,&n<0\thinspace,\enspace n=\text{odd}\\ -1\thinspace,&n<0\thinspace,\enspace n=\text{even}\end{cases}$ (Eq. 3.6.1)
@@
$(-1)^{n}=\begin{cases}-1\thinspace,&n=\text{odd}\\ \thinspace 1\thinspace,&n=\text{even}\end{cases}$
@@
$(-1)^{n-1}=\begin{cases}\thinspace 1\thinspace,&n=\text{odd}\\ -1\thinspace,&n=\text{even}\end{cases}$
Let’s now address the case of positive inputs. Since we raise the base $(-1)$ to some power, if the base becomes $(+1)$ for $n>0$, then the result will always be 1 regardless of the power. That is true because $1^{n}=1,\enspace n\in\mathbb{R}\thinspace$. To ensure that behavior we can use $\theta(n)$ for the base:
$\eta(n)=\theta(n)^{n-1}\thinspace,\quad n\in\mathbb{Z}\thinspace.$ (Eq. 3.6.2)
@@
$\eta(n)=\theta(n)^{\lceil n\rceil-1}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.6.3)
@@
$(-1)^{\lceil-5\rceil-1}=(-1)^{-\lceil-5\rceil-1}=1\thinspace.$
@@
$\eta(n)=\theta(n)^{-\lceil n\rceil-1}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.6.5)
@@ 3.7 Synopsis
Refer to Eq. 3.7.1 for a reminder of the output patterns when these functions are utilized in Section 4. Additional F.F. will be introduced in Section 5, to be employed in Section 6.
4 Recursive Roman factorial
@@ 4.1 Introduction
$\lfloor n\rceil!=\begin{cases}n!&,\enspace n\in\mathbb{N}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 4.1.1)
@@
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 4.1.2)
Here, $\mathbb{N}_{0}$ refers to the set of natural numbers plus 0, while $\mathbb{Z}^{-}$ represents negative integers. More about number sets in Addendum 9.1.
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace,\end{cases}$ (Eq. 4.1.3)
@@
[…]
We will call the generalization of Eq. 4.1.3 as first generalization, and that of Eq. 4.1.1 as second generalization. This is because we will address the easier one in the beginning and the more complicated one directly afterwards.
@@
Despite the rewriting, the resulting relationships will still be based on a recursive factorial. In Section 6, we will develop a new factorial definition from scratch.
4.2 First generalization step 1: $\theta(n)$
In Subsections 4.2 and 4.3 we aim to combine the two instances of the doubly-recursive definition of the Roman factorial:
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace,\end{cases}$ (Eq. 4.2.1)
@@ 4.2 First generalization step 1: $\theta(n)$
$\begin{cases}n\cdot\lfloor n-1\rceil!\\[3.68027pt] \cfrac{1}{n+1}\cdot\lfloor n+1\rceil!\end{cases}=\begin{cases}n^{+1}\cdot\lfloor n-(+1)\rceil!\\[3.68027pt] (n+1)^{-1}\cdot\lfloor n-(-1)\rceil!\end{cases}$ (Eq. 4.2.3)
In this way, we can see that there are two occurrences of a function in each case. That function is $+1$ when $n\in\mathbb{N}$ and $-1$ when $n\in\mathbb{Z}^{-}\setminus\{-1\}$ ($n$ is a negative integer except $-1$).
@@
$\theta(n)=\cfrac{\delta(n)}{|\delta(n)|}=\begin{cases}1&,\enspace n\in\mathbb{N}\\ -1&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}.\end{cases}$ (Eq. 4.2.4)
@@
$\lfloor n\rceil!=\begin{cases}n^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] (n+1)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace,\end{cases}$ (Eq. 4.2.6)
@@
[…]
4.3 First generalization step 2: $\xi^{\prime}(n)$
We observe that in the Eq. 4.2.5, the only difference between the cases is the addition of 1 in the first term, which is absent when $n\in\mathbb{N}$.
@@ 4.3 First generalization step 2: $\xi^{\prime}(n)$
$\lfloor n\rceil!=\begin{cases}(n+0)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] (n+1)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace.\end{cases}$ (Eq. 4.3.1)
We notice that this additive term is 0 when $n\in\mathbb{N}$ and 1 when $n\in\mathbb{Z}^{-}\setminus\{-1\}$. This behavior is described by the F.F. $\xi^{\prime}(n)$:
$\xi^{\prime}(n)=\begin{cases}0&,\enspace n\in\mathbb{N}\\ 1&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace.\end{cases}$ (Eq. 4.3.2)
Incorporating this foundational function into Eq. 4.3.1 results in the following formulation:
$\begin{cases}\big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] \big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\end{cases}$ (Eq. 4.3.3)
@@
$\lfloor n\rceil!=\big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!\thinspace,\enspace n\in\mathbb{Z}\setminus\{0,\enspace-1\}\thinspace,$ (Eq. 4.3.4)
@@
[…]
This result concludes the first generalization, so in the next subsections we will address the second one.
4.4 Second generalization step 1: $\eta(n)$
@@
$\lfloor n\rceil!=\begin{cases}n!&,\enspace n\in\mathbb{N}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 4.4.1)
@@
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 4.4.2)
@@
$\lfloor n\rceil!=\begin{cases}1\cdot n!&,\enspace n\in\mathbb{N}_{0}\\ (-1)^{-n-1}\cdot\cfrac{1}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.4.3)
@@
$\eta(n)=\theta(n)^{-n-1}=\begin{cases}1&,\enspace n\in\mathbb{N}_{0}\\ (-1)^{-n-1}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.4.4)
@@
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot n!&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot\cfrac{1}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.4.5)
@@
4.5 Second generalization step 2: $\theta(n)$

### v7

@@ 3.3 The function $\theta(n)$
$\theta(n)=\begin{cases}+1\,,\quad n\geq 0\\ -1\,,\quad n<0\end{cases},\enspace n\in\mathbb{R}\,.$ (Eq. 3.3.1)
@@
$\theta(n)=\frac{\delta(n)}{\left|\delta(n)\right|}\enspace,\quad n\in\mathbb{R}\,.$ (Eq. 3.3.2)
@@ 3.4 The function $\xi(n)$
$\xi(n)=\begin{cases}1\,,\quad n\geq 0\\ 0\,,\quad n<0\end{cases},\enspace n\in\mathbb{R}\,.$ (Eq. 3.4.1)
Let’s construct $\xi(n)$ using $\theta(n)$ step by step. To begin with, we note that the output pattern of $\theta(n)$ is $[-1,1,1]$. Adding 1 to $\theta(n)$ essentially adds to each output, obtaining $[0,2,2]$. Dividing $\big{(}1+\theta(n)\big{)}$ by 2 yields the desired pattern $[0,1,1]$. Thus, we arrive at the universal definition of $\xi(n)$:
$\xi(n)=\frac{1+\theta(n)}{2}\>,\quad n\in\mathbb{R}\,.$ (Eq. 3.4.2)
@@
This function has similarities to true/false states in Boolean logic, which is commonly used in programming. Specifically, the expressions $\xi(n\geq 0)=1$ and $\xi(n<0)=0$ represent the true/false statements of the condition $n\geq 0$. Furthermore, it’s noteworthy that the most subsequent F.F. can also be expressed in Boolean algebra [^11] More about Boolean algebra in Addendum 9.5..
@@ 3.5 The function $\xi^{\prime}(n)$
The function $\xi^{\prime}(n)$, marked with (), bears resemblance to $\xi(n)$, although it is not its derivative. In this paper, the () symbol signifies similarity between functions rather than indicating a derivative. Notably, the F.F. are devoid of derivatives since they are constructed upon the floor function, which itself lacks a derivative at integers. Derivatives will not be discussed within the scope of this work [^12] It should be mentioned again that the names of the F.F. are arbitrarily assigned Greek or Latin letters and have no relation to other functions with the same name. The naming process was conducted without any specific goal or consideration of potential overlaps with pre-existing functions..
@@
$\xi^{\prime}(n)=\frac{1-\theta(n)}{2}\>,\quad n\in\mathbb{R}\,.$ (Eq. 3.5.1)
@@
$\xi^{\prime}(n)=1-\xi(n)\,,\quad n\in\mathbb{R}\,.$ (Eq. 3.5.2)
In Boolean algebra, $\xi^{\prime}(n)$ yields the true/false states of the condition $n<0\,$, thereby exhibiting an output pattern of $[1,0,0]$.
@@ 3.6 The function $\eta(n)$
$\eta(n)=\begin{cases}\,1\,,&n\geq 0\\ \,1\,,&n<0\,,\enspace n=\text{odd}\\ -1\,,&n<0\,,\enspace n=\text{even}\,.\end{cases}$ (Eq. 3.6.1)
@@
$(-1)^{n}=\begin{cases}-1\,,&n=\text{odd}\\ \,1\,,&n=\text{even}\,.\end{cases}$
@@
$(-1)^{n-1}=\begin{cases}\,1\,,&n=\text{odd}\\ -1\,,&n=\text{even}\,.\end{cases}$
Let’s now address the case of positive inputs. Since we raise the base $(-1)$ to some power, if the base becomes $(+1)$ for $n>0$, then the result will always be 1 regardless of the power. That is true because $1^{n}=1,\enspace n\in\mathbb{R}\,$. To ensure that behavior we can use $\theta(n)$ for the base:
$\eta(n)=\theta(n)^{n-1}\,,\quad n\in\mathbb{Z}\,.$ (Eq. 3.6.2)
@@
$\eta(n)=\theta(n)^{\lceil n\rceil-1}\,,\quad n\in\mathbb{R}\,.$ (Eq. 3.6.3)
@@
$(-1)^{\lceil-5\rceil-1}=(-1)^{-\lceil-5\rceil-1}=1\,.$
@@
$\eta(n)=\theta(n)^{-\lceil n\rceil-1}\,,\quad n\in\mathbb{R}\,.$ (Eq. 3.6.5)
@@ 3.7 Synopsis
$\delta(n)$ $=\lfloor n\rfloor+0.5$ [,+,+[-,+,+][ - , + , + ] (Eq. 3.2.3)
$\theta(n)$ $=\cfrac{\delta(n)}{\left|\delta(n)\right|}$ $[-1,1,1]$ (Eq. 3.3.2)
$\xi(n)$ $=\cfrac{1+\theta(n)}{2}$ $[0,1,1]$ (Eq. 3.4.2)
$\xi^{\prime}(n)$ $=\cfrac{1-\theta(n)}{2}$ $[1,0,0]$ (Eq. 3.5.1)
@@
You can refer to these equations for a reminder of the output patterns when these functions are utilized in Section 4. Additional F.F. will be introduced in Section 5, to be employed in Section 6.
4 Recursive generalizations of $\lfloor n\rceil!$
@@ 4.1 Introduction
$\lfloor n\rceil!=\begin{cases}n!&,\enspace n\in\mathbb{Z}^{+}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 2.2.1)
@@
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
Here, $\mathbb{Z}^{+}_{0}$ refers to the set of natural numbers plus 0, while $\mathbb{Z}^{-}$ represents negative integers. More about number sets in Addendum 9.1.
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 2.2.3)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
We will call the generalization of Eq. 2.2.3 as a recursive definition, and that of Eq. 2.2.1 as a Roman-like definition. This is because we will address the easier one in the beginning and the more complicated one directly afterwards.
@@
Despite the rewriting, the resulting relationships will still be based on a recursive factorial. Later, in Section 6, we will develop a new factorial definition from scratch.
4.2 Recursive definition step 1: $\theta(n)$
In Subsections 4.2 and 4.3 (the rest of this page) we aim to combine the two instances of the doubly-recursive definition of the Roman factorial:
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 2.2.3)
@@ 4.2 Recursive definition step 1: $\theta(n)$
$\begin{cases}n\cdot\lfloor n-1\rceil!\\[3.68027pt] \cfrac{1}{n+1}\cdot\lfloor n+1\rceil!\end{cases}=\begin{cases}n^{+1}\cdot\lfloor n-(+1)\rceil!\\[3.68027pt] (n+1)^{-1}\cdot\lfloor n-(-1)\rceil!\end{cases}$
In this way, we can see that there are two occurrences of a function in each case. That function is $+1$ when $n\in\mathbb{Z}^{+}$ and $-1$ when $n\in\mathbb{Z}^{-}\setminus\{-1\}$ (meaning $n$ is a negative integer except $-1$).
@@
$\theta(n)=\cfrac{\delta(n)}{|\delta(n)|}=\begin{cases}1&,\enspace n\in\mathbb{Z}^{+}\\ -1&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,.\end{cases}$ (Eq. 3.3.2)
@@
$\lfloor n\rceil!=\begin{cases}n^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] (n+1)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 4.2.2)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
4.3 Recursive definition step 2: $\xi^{\prime}(n)$
We observe that in the Eq. 4.2.1, the only difference between the cases is the addition of 1 in the first term, which is absent when $n\in\mathbb{Z}^{+}$.
@@ 4.3 Recursive definition step 2: $\xi^{\prime}(n)$
$\lfloor n\rceil!=\begin{cases}(n+0)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] (n+1)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,.\end{cases}$ (Eq. 4.3.1)
Notice that this additive term is 0 when $n\in\mathbb{Z}^{+}$ and 1 when $n\in\mathbb{Z}^{-}\setminus\{-1\}$. This behavior is described by the F.F. $\xi^{\prime}(n)$:
$\xi^{\prime}(n)=\begin{cases}0&,\enspace n\in\mathbb{Z}^{+}\\ 1&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,.\end{cases}$ (Eq. 3.5.1)
Incorporating this F.F. into Eq. 4.3.1 results in the following formulation:
$\begin{cases}\big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\end{cases}$
@@
$\lfloor n\rceil!=\big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!\,,\quad n\in\mathbb{Z}\setminus\{0,\enspace-1\}\,,$ (Eq. 4.3.2)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
This result concludes the recursive generalization, so in the next subsections we will address the Roman-like one.
4.4 Roman-like definition step 1: $\eta(n)$
@@
$\lfloor n\rceil!=\begin{cases}n!&,\enspace n\in\mathbb{Z}^{+}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 2.2.1)
@@
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
@@
$\lfloor n\rceil!=\begin{cases}1\cdot n!&,\enspace n\in\mathbb{Z}^{+}_{0}\\ (-1)^{-n-1}\cdot\cfrac{1}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.4.1)
@@
$\eta(n)=\theta(n)^{-n-1}=\begin{cases}1&,\enspace n\in\mathbb{Z}^{+}_{0}\\ (-1)^{-n-1}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 3.6.5)
@@
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot n!&,\enspace n\in\mathbb{Z}^{+}_{0}\\ \eta(n)\cdot\cfrac{1}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.4.2)
@@
4.5 Roman-like definition step 2: $\theta(n)$